# Holmes: Towards Effective and Harmless Model Ownership Verification to Personalized Large Vision Models via Decoupling Common Features

### v1

@@ V-C Effects of Key Hyper-parameters
V-C1 Effects of Modification Rate
TABLE VI: Effects of modification rate of the poisoned shadow model.
Model Stealing $\rightarrow$ Direct-copy Fine-tuning Distillation DFKD Label-query Logit-query Independent
@@ V-C1 Effects of Modification Rate
0% -0.03 0.40 -0.21 0.99 -0.09 0.12 -0.01 0.12 $-10^{-3}$ 0.33 -0.30 0.51 0.07 $10^{-4}$
5% 0.17 $10^{-11}$ 0.02 0.01 0.16 $10^{-8}$ 0.23 $10^{-9}$ 0.05 $10^{-4}$ 0.27 $10^{-16}$0.04 0.30
10% 0.18 $10^{-14}$ 0.12 $10^{-9}$ 0.18 $10^{-12}$ 0.13 $10^{-9}$ 0.05 $10^{-4}$ 0.14 $10^{-8}$0.03 0.20
15% 0.19 $10^{-12}$ $10^{-3}$ $10^{-3}$ 0.15 $10^{-8}$ 0.23 $10^{-16}$ 0.03 $10^{-3}$ 0.22 $10^{-21}$0.20 0.29
20% 0.19 $10^{-15}$ 0.14 $10^{-5}$ 0.07 $10^{-10}$ 0.12 $10^{-3}$ 0.02 $10^{-3}$ 0.09 $10^{-7}$0.01 0.04
TABLE VII: Effects of drop rate of the benign shadow model.
Model Stealing $\rightarrow$ Direct-copy Fine-tuning Distillation DFKD Label-query Logit-query Independent
@@
0% 0.29 $10^{-12}$0.07 0.38 $10^{-5}$ 0.59 -0.01 0.63 -0.02 0.02 -0.04 0.22 0.13 $10^{-7}$
10% 0.20 $10^{-14}$ 0.01 $10^{-9}$ 0.17 $10^{-9}$ 0.17 $10^{-13}$ 0.04 $10^{-3}$ 0.08 $10^{-5}$0.03 0.15
30% 0.15 $10^{-14}$ 0.12 $10^{-9}$ 0.26 $10^{-11}$ 0.25 $10^{-14}$ 0.03 $10^{-4}$ 0.15 $10^{-8}$0.07 0.19
50% 0.18 $10^{-14}$ 0.12 $10^{-9}$ 0.18 $10^{-12}$ 0.13 $10^{-9}$ 0.05 $10^{-4}$ 0.14 $10^{-8}$0.03 0.20
70% 0.21 $10^{-26}$ 0.16 $10^{-7}$ 0.30 $10^{-18}$ 0.25 $10^{-25}$ 0.20 $10^{-11}$ 0.25 $10^{-15}$0.03 0.19
TABLE VIII: Effects of number of verification samples.
@@
$m=50$ 0.17 $10^{-7}$ 0.11 $10^{-9}$ 0.18 $10^{-7}$ 0.13 $10^{-4}$ 0.05 $10^{-3}$ 0.12 $10^{-6}$0.02 0.14
$m=100$ 0.18 $10^{-14}$ 0.12 $10^{-9}$ 0.18 $10^{-12}$ 0.13 $10^{-9}$ 0.05 $10^{-4}$ 0.14 $10^{-8}$0.03 0.20
$m=150$ 0.18 $10^{-14}$ 0.12 $10^{-9}$ 0.18 $10^{-15}$ 0.14 $10^{-9}$ 0.05 $10^{-4}$ 0.14 $10^{-10}$0.03 0.28
$m=200$ 0.19 $10^{-19}$ 0.12 $10^{-9}$ 0.17 $10^{-15}$ 0.14 $10^{-10}$ 0.05 $10^{-4}$ 0.14 $10^{-10}$0.03 0.36
TABLE IX: Effects of model structures.
@@
$\Delta\mu$ p-value $\Delta\mu$ p-value $\Delta\mu$ p-value $\Delta\mu$ p-value
Direct-copy 0.24 $10^{-16}$ 0.28 $10^{-31}$ 0.11 $10^{-6}$ 0.18 $10^{-14}$
Fine-tuning 0.14 $10^{-3}$ 0.28 $10^{-21}$ 0.01 $10^{-3}$ 0.12 $10^{-9}$
Distillation 0.13 $10^{-4}$ 0.22 $10^{-20}$ 0.06 $10^{-6}$ 0.18 $10^{-12}$
DFKD 0.01 $10^{-3}$ 0.23 $10^{-10}$ 0.02 $10^{-3}$ 0.13 $10^{-9}$
Label-query 0.13 $10^{-3}$ 0.01 $10^{-3}$ 0.03 $10^{-3}$ 0.05 $10^{-4}$
Logit-query 0.17 $10^{-3}$ 0.24 $10^{-24}$ 0.06 $10^{-3}$ 0.14 $10^{-8}$
Independent -0.33 0.99 -0.05 0.29 -0.08 0.60 -0.03 0.20
In general, a higher modification rate $\gamma\%$ leads to greater disruption of dataset-specific features and less preservation of common features in the poisoned shadow model. As shown in Table VI, when the modification rate is 0% ($i.e.$, no poisoning is applied), the method fails to distinguish between stolen and benign models, leading to misjudgments. This is because the output difference between the victim model and the unpoisoned shadow model (which is essentially similar to the victim model) fails to isolate the dataset-specific features, causing the meta-classifier to fail to identify dataset-specific features. Conversely, when $\gamma\%\in\{5\%,10\%,15\%,20\%\}$, the method achieves accurate verification across all cases. Specifically, for the direct-copy attack, a $\gamma\%$ of 5% yields a p-value as low as $10^{-11}$, indicating effective verification results. As $\gamma\%$ increases from 5% to 20%, the p-values exhibit a slight downward trend ($e.g.$, from $10^{-11}$ to $10^{-15}$ for direct-copy attacks), reflecting the reliable disruption of dataset-specific features.
V-C2 Effects of Drop Rate
For the benign shadow model, we drop $\lambda\%$ of the most overfitted samples from the victim model’s training dataset to ensure that the benign shadow model learns different dataset-specific features while preserving similar common features. This approach allows the output difference between the benign shadow model and the poisoned shadow model to represent similar but distinct dataset-specific features, enabling the meta-classifier to better distinguish between independent benign models fine-tuned on similar datasets and actual stolen models. In general, a higher drop rate leads to greater differences in dataset-specific features but reduces the preservation of common features in the benign shadow model. As shown in Table VII, when $\lambda\%=0\%$, our method fails to provide ownership verification in almost all cases because the benign shadow model learns dataset-specific features that are too similar to those of the victim model. For drop rates of $\lambda\%\in\{10\%,30\%,50\%,70\%\}$, our method succeeds in providing effective verification in all cases.
V-C3 Effects of the Number of Verification Samples
Our method requires specifying the number of verification samples from $\mathcal{D}_{s}$ ($i.e.$, $m$), used in the hypothesis test process. Generally, increasing $m$ mitigates the randomness inherent in the sampling process, thereby enhancing confidence of the verification results. This is probably the main reason why the p-value consistently decreases as $m$ increases. For example, as shown in Table VIII, the p-value decreases from $10^{-6}$ ($m=50$) to $10^{-10}$ ($m=200$) for the logit-query attack. However, larger $m$ introduces higher computational costs due to increased model queries and processing overhead. This creates a trade-off between verification confidence and computational efficiency, requiring defenders to balance $m$ based on specific deployment constraints and verification requirements.
V-C4 Effects of Model Structures
In this section, we discuss the generalization of our method across different model architectures. Specifically, we consider two cases: (1) the stolen models have the same structures as the victim model, and (2) the stolen models have different structures from the victim model. In the first case, we conduct experiments on BEiT to BEiT. In the second case, we consider three settings, including ResNet34 to ResNet18, VGG19 to VGG16, and CLIP to ResNet18. The fine-tuning process of victim models and the model stealing settings remain consistent with those described in Section V-A. As shown in Table IX, our method remains effective under all settings without misjudgments.
V-D Ablation Study
There are three key parts contained in our method, including and (1) the poisoned shadow model $P$, (2) the benign shadow model $B$, and (3) the selection strategy for $\mathcal{D}_{s}$ and $\mathcal{D}_{f}$. In this section, we evaluate the effectiveness of each component.
V-D1 The Effectiveness of the Poisoned Shadow Model
TABLE X: The Effectiveness of shadow models.
@@ V-D1 The Effectiveness of the Poisoned Shadow Model
$\Delta\mu$ p-value $\Delta\mu$ p-value $\Delta\mu$ p-value
Direct-copy 0.18 $10^{-14}$ 0.32 $10^{-32}$ 0.47 $10^{-45}$
Fine-tuning 0.12 $10^{-9}$ 0.25 $10^{-30}$ 0.39 $10^{-7}$
Distillation 0.18 $10^{-12}$ 0.32 $10^{-30}$ 0.32 $10^{-3}$
DFKD 0.13 $10^{-9}$ 0.36 $10^{-36}$ 0.30 0.01
Label-query 0.05 $10^{-4}$0.18 0.19 0.10 $10^{-3}$
Logit-query 0.14 $10^{-8}$0.21 0.22 0.08 0.01
Independent -0.03 0.20 0.20 $10^{-15}$ 0.25 $10^{-10}$
@@
Similar to the verification process described in Section IV-C, we also adopt hypothesis test to mitigate the randomness. The hypothesis test is as follows:
@@
Let $\bm{X}$ be a random variable representing samples from $\mathcal{D}_{s}$. Define $\mu_{S}$ and $\mu_{B}$ as the posterior probabilities of the events $C(S(\bm{x}))=+1$ and $C(S(\bm{x}))=-1$, respectively. Given the null hypothesis $H_{0}:\mu_{B}+\tau=\mu_{S}$ ($H_{1}:\mu_{B}+\tau<\mu_{S}$) where $\tau\in[0,1]$, we claim that the suspicious model $S$ is stolen from the victim (with $\tau$certainty) if and only if $H_{0}$ is rejected.
As shown in Table X, the absence of a poisoned shadow model not only leads to the false positive of the independent model but also causes misjudgment of label-query attack and logit-query attack. This is likely because the verification process fails to capture dataset-specific features. Without a poisoned shadow model to decouple these common features, the meta-classifier becomes unable to distinguish between independent models and actual stolen models, resulting in model ownership verification failures.
V-D2 The Effectiveness of the Benign Shadow Model
$C$ is trained only on positive samples using the Isolation Forest method [62]. To mitigate the randomness in the model ownership verification process, we also employ a hypothesis test approach, formulated as follows:
Definition 4.
Let $\bm{X}$ be a random variable representing samples from $\mathcal{D}_{s}$. Define $\mu_{S}$ and $\mu_{B}$ as the posterior probabilities of the events $C(S(\bm{x})-P(\bm{x}))=+1$ and $C(S(\bm{x})-P(\bm{x}))=-1$, respectively. Given the null hypothesis $H_{0}:\mu_{B}+\tau=\mu_{S}$ ($H_{1}:\mu_{B}+\tau<\mu_{S}$) where $\tau\in[0,1]$, we claim that the suspicious model $S$ is stolen from the victim (with $\tau$certainty) if and only if $H_{0}$ is rejected.
As shown in Table X, without the benign shadow model, the independent model is misjudged as a stolen model. This is probably because the independent model contains dataset-specific features similar to those of the victim model. Without being able to distinguish the victim model’s dataset-specific features from the similar but distinct dataset-specific features of other models, the meta-classifier fails to distinguish independent models from actual stolen models.
V-D3 The Effectiveness of the Selection Strategy for $\mathcal{D}_{s}$ and $\mathcal{D}_{f}$
TABLE XI: Effects of $\mathcal{D}_{s}$ selection.
@@ V-D3 The Effectiveness of the Selection Strategy for $\mathcal{D}_{s}$ and $\mathcal{D}_{f}$
Fine-tuning 0.12 $\mathbf{10^{-9}}$0.04 0.37 0.03 $10^{-4}$
Distillation 0.18 $\mathbf{10^{-12}}$0.04 0.43 0.10 $10^{-4}$
DFKD 0.13 $\mathbf{10^{-9}}$ 0.13 $10^{-7}$ 0.07 $10^{-5}$
Label-query 0.05 $\mathbf{10^{-4}}$ 0.04 $\mathbf{10^{-4}}$ 0.04 $\mathbf{10^{-4}}$
Logit-query 0.14 $10^{-8}$ 0.02 $\mathbf{10^{-11}}$ 0.20 $10^{-5}$
Independent -0.03 0.20 0.18 $10^{-13}$0.05 0.59
TABLE XII: Effects of $\mathcal{D}_{f}$ selection.
$\mathcal{D}_{s}\rightarrow$ Lowest Loss Highest Loss Random

### v2

@@ 5.3 Effects of Hyper-parameters
5.3.1 Effects of Modification Rate
Table 6: Effects of modification rate of the poisoned shadow model.
Model Stealing $\rightarrow$ Direct-copy Fine-tuning Data Distill Data-free Distill Hard Distill Soft Distill Independent
@@ 5.3.1 Effects of Modification Rate
0% -0.03 0.40 -0.21 0.99 -0.09 0.12 -0.01 0.12 $10^{-3}$ 0.33 -0.30 0.51 0.07 $10^{-4}$
5% 0.17 $10^{-11}$ 0.02 0.01 0.16 $10^{-8}$ 0.23 $10^{-9}$ 0.05 $10^{-4}$ 0.27 $10^{-16}$ -0.04 0.30
10% 0.18 $10^{-14}$ 0.12 $10^{-9}$ 0.18 $10^{-12}$ 0.13 $10^{-9}$ 0.05 $10^{-4}$ 0.14 $10^{-8}$ -0.03 0.20
15% 0.19 $10^{-12}$ $10^{-3}$ $10^{-3}$ 0.15 $10^{-8}$ 0.23 $10^{-16}$ 0.03 $10^{-3}$ 0.22 $10^{-21}$ -0.20 0.29
20% 0.19 $10^{-15}$ 0.14 $10^{-5}$ 0.07 $10^{-10}$ 0.12 $10^{-3}$ 0.02 $10^{-3}$ 0.09 $10^{-7}$ -0.01 0.04
Table 7: Effects of drop rate of the benign shadow model.
Model Stealing $\rightarrow$ Direct-copy Fine-tuning Data Distill Data-free Distill Hard Distill Soft Distill Independent
@@
0% 0.29 $10^{-12}$ -0.07 0.38 $10^{-5}$ 0.59 -0.01 0.63 -0.02 0.02 -0.04 0.22 0.13 $10^{-7}$
[…]
30% 0.15 $10^{-14}$ 0.12 $10^{-9}$ 0.26 $10^{-11}$ 0.25 $10^{-14}$ 0.03 $10^{-4}$ 0.15 $10^{-8}$ -0.07 0.19
50% 0.18 $10^{-14}$ 0.12 $10^{-9}$ 0.18 $10^{-12}$ 0.13 $10^{-9}$ 0.05 $10^{-4}$ 0.14 $10^{-8}$ -0.03 0.20
70% 0.21 $10^{-26}$ 0.16 $10^{-7}$ 0.30 $10^{-18}$ 0.25 $10^{-25}$ 0.20 $10^{-11}$ 0.25 $10^{-15}$ -0.03 0.19
Table 8: Effects of number of verification samples.
@@
$m=50$ 0.17 $10^{-7}$ 0.11 $10^{-9}$ 0.18 $10^{-7}$ 0.13 $10^{-4}$ 0.05 $10^{-3}$ 0.12 $10^{-6}$ -0.02 0.14
$m=100$ 0.18 $10^{-14}$ 0.12 $10^{-9}$ 0.18 $10^{-12}$ 0.13 $10^{-9}$ 0.05 $10^{-4}$ 0.14 $10^{-8}$ -0.03 0.20
$m=150$ 0.18 $10^{-14}$ 0.12 $10^{-9}$ 0.18 $10^{-15}$ 0.14 $10^{-9}$ 0.05 $10^{-4}$ 0.14 $10^{-10}$ -0.03 0.28
$m=200$ 0.19 $10^{-19}$ 0.12 $10^{-9}$ 0.17 $10^{-15}$ 0.14 $10^{-10}$ 0.05 $10^{-4}$ 0.14 $10^{-10}$ -0.03 0.36
In general, a higher modification rate $\gamma\%$ leads to greater disruption of dataset-specific features and less preservation of common features in model $P$. As shown in Table 6, when the modification rate is 0% ($i.e.$, no poisoning is applied), the method fails to distinguish between stolen and benign models, leading to misjudgments. This is because the output difference between the victim model and the unpoisoned shadow model (which is essentially similar to the victim model) fails to isolate the dataset-specific features, causing the meta-classifier to fail to identify dataset-specific features. Conversely, when $\gamma\%\in\{5\%,10\%,15\%,20\%\}$, the method achieves accurate verification across all cases. Specifically, for the Direct-copy attack, $\gamma\%$ of 5% yields a p-value as low as $10^{-11}$, indicating effective verification. As $\gamma\%$ increases from 5% to 20%, the p-values exhibit a slight downward trend ($e.g.$, from $10^{-11}$ to $10^{-15}$ for Direct-copy attacks), reflecting reliable disruption of dataset-specific features.
5.3.2 Effects of Drop Rate
We drop $\lambda\%$ training samples with the lowest loss values from $V$’s training dataset to ensure that $B$ learns different dataset-specific features while preserving similar common features. This approach allows the output difference between $B$ and $P$ to represent similar but distinct dataset-specific features. This enables the meta-classifier to better distinguish between independent benign models fine-tuned on similar datasets and actual stolen models. In general, a higher drop rate leads to greater differences in dataset-specific features but reduces the preservation of common features. As shown in Table 7, when $\lambda\%=0\%$, our method fails to provide ownership verification in almost all cases because the benign shadow model learns dataset-specific features that are too similar to those of the victim model. For drop rates of $\lambda\%\in\{10\%,30\%,50\%,70\%\}$, our method succeeds in providing effective verification in all cases.
Table 9: Effects of model structures.
@@ 5.3.2 Effects of Drop Rate
[…]
CLIP 0.18 $10^{-14}$ 0.12 $10^{-9}$ 0.18 $10^{-12}$ 0.13 $10^{-9}$} 0.05 $10^{-4}$} 0.14 $10^{-8}$} -0.03 0.20
5.3.3 Effects of the Number of Verification Samples
Holmes requires specifying the number of verification samples from $\mathcal{D}_{s}$ ($i.e.$, $m$) used in the hypothesis test. Generally, increasing $m$ mitigates the randomness inherent in the sampling process, thereby enhancing the confidence of the verification results. This is probably the main reason why the p-value consistently decreases as $m$ increases. For example, as shown in Table 8, the p-value decreases from $10^{-6}$ ($m=50$) to $10^{-10}$ ($m=200$) for the soft distillation attack. However, larger $m$ introduces higher computational costs due to increased model queries and processing overhead. This creates a trade-off between verification confidence and computational efficiency, requiring defenders to balance $m$ based on specific deployment constraints and verification requirements.
5.3.4 Effects of Model Structures
In this section, we discuss the generalization of our method across different model architectures. Specifically, we consider two cases: (1) the stolen models have the same structures as the victim model, and (2) the stolen models have different structures from the victim model. In the first case, we conduct experiments on BEiT to BEiT. In the second case, we consider three settings, including ResNet34 to ResNet18, VGG19 to VGG16, and CLIP to ResNet18. The fine-tuning process of victim models and the model stealing settings remain consistent with those described in Section 5.1. As shown in Table 9, our method remains effective under all settings without misjudgments.
5.4 Ablation Study
There are three key parts contained in our method, including (1) the poisoned shadow model $P$, (2) the benign shadow model $B$, and (3) the selection strategy for $\mathcal{D}_{s}$ and $\mathcal{D}_{f}$. In this section, we evaluate the effectiveness of each component.
5.4.1 The Effectiveness of the Poisoned Shadow Model
@@
Similar to the verification process described in Section 4.3, we also adopt hypothesis test to mitigate the randomness as follows:
@@
Let $\bm{X}$ be a random variable representing samples from $\mathcal{D}_{s}$. Define $\mu_{S}$ and $\mu_{B}$ as the posterior probabilities of the events $C(S(\bm{x}))=+1$ and $C(S(\bm{x}))=-1$, respectively. Given the null hypothesis $H_{0}:\mu_{B}+\tau=\mu_{S}$ ($H_{1}:\mu_{B}+\tau<\mu_{S}$) where $\tau\in[0,1]$, we claim that the suspicious model $S$ is stolen from the victim (with $\tau$-certainty) if and only if $H_{0}$ is rejected.
Table 10: The effectiveness of shadow models.
[…]
$\Delta\mu$ p-value $\Delta\mu$ p-value $\Delta\mu$ p-value
Direct-copy 0.18 $10^{-14}$ 0.32 $10^{-32}$ 0.28 $10^{-24}$
Fine-tuning 0.12 $10^{-9}$ 0.25 $10^{-30}$ 0.23 $10^{-16}$
Data Distill 0.18 $10^{-12}$ 0.32 $10^{-30}$ 0.11 $10^{-6}$
Data-free Distill 0.13 $10^{-9}$ 0.36 $10^{-36}$ 0.14 $10^{-4}$
Hard Distill 0.05 $10^{-4}$ -0.18 0.19 0.02 $10^{-3}$
Soft Distill 0.14 $10^{-8}$ -0.21 0.22 0.07 $10^{-3}$
Independent -0.03 0.20 0.11 $10^{-3}$ 0.07 $10^{-7}$
As shown in Table 10, the absence of a poisoned shadow model not only leads to the false positive of the independent model but also causes misjudgment of hard distillation and soft distillation attacks. This is likely because the verification process fails to capture dataset-specific features. Without a poisoned shadow model to decouple these common features, the meta-classifier becomes unable to distinguish between independent models and actual stolen models, resulting in model ownership verification failures.
Table 11: Effects of $\mathcal{D}_{s}$ selection.
@@
Fine-tuning 0.12 $\mathbf{10^{-9}}$ -0.04 0.37 0.03 $10^{-4}$
Data Distill 0.18 $\mathbf{10^{-12}}$ -0.04 0.43 0.10 $10^{-4}$
Data-free Distill 0.13 $\mathbf{10^{-9}}$ 0.13 $10^{-7}$ 0.07 $10^{-5}$
Hard Distill 0.05 $\mathbf{10^{-4}}$ 0.04 $\mathbf{10^{-4}}$ 0.04 $\mathbf{10^{-4}}$
Soft Distill 0.14 $10^{-8}$ 0.02 $\mathbf{10^{-11}}$ 0.20 $10^{-5}$
Independent -0.03 0.20 0.18 $10^{-13}$ -0.05 0.59
5.4.2 The Effectiveness of the Benign Shadow Model
$C$ is trained only on positive samples using the Isolation Forest method [liu2008isolation]. To mitigate the randomness in the model ownership verification process, we also employ a hypothesis test approach, formulated as follows:
Definition 4.
Let $\bm{X}$ be a random variable representing samples from $\mathcal{D}_{s}$. Define $\mu_{S}$ and $\mu_{B}$ as the probabilities of the events $C(S(\bm{x})-P(\bm{x}))=+1$ and $C(S(\bm{x})-P(\bm{x}))=-1$, respectively. These probabilities are estimated by the frequencies of +1 and -1 outputs of $C$ over $z$ randomly selected samples from $\mathcal{D}_{s}$. Given the null hypothesis $H_{0}:\mu_{B}+\tau=\mu_{S}$ ($H_{1}:\mu_{B}+\tau<\mu_{S}$) where $\tau\in[0,1]$, we claim that the suspicious model $S$ is stolen from the victim (with $\tau$-certainty) if and only if $H_{0}$ is rejected.
As shown in Table 10, when $z=10,m=100$, without the benign shadow model, the independent model is misjudged as a stolen model. This is probably because the independent model shares similar dataset-specific features with the victim model. Unable to distinguish the victim’s dataset-specific features from the analogous yet distinct ones of other models, the classifier confuses independent models with actual stolen ones.
5.4.3 The Effectiveness of the Selection Strategy for $\mathcal{D}_{s}$ and $\mathcal{D}_{f}$
Our method selects $\gamma\%$ of samples with the lowest loss values to form $\mathcal{D}_{s}$ and drops $\lambda\%$ with the lowest loss values to form $\mathcal{D}_{f}$, as these samples likely correspond to the model’s propensity for memorizing dataset-specific patterns rather than common features [ilyas2019adversarial]. To verify the effectiveness of this selection strategy, we conduct experiments using two alternative approaches: selecting the $\gamma\%$ ($\lambda\%$) of samples with the highest loss values and randomly selecting $\gamma\%$ ($\lambda\%$) of the samples.
As shown in Table 11-12, selecting samples with the lowest loss values yields the best verification results in almost all cases for both $\mathcal{D}_{s}$ and $\mathcal{D}_{f}$ selections. For $\mathcal{D}_{s}$, selecting samples with the highest loss values fails in cases of fine-tuning attacks, distillation attacks, and identifying the independent model. This failure is probably because the samples with the highest loss values are less effectively learned by the victim model compared to other samples, making them inadequate to accurately capture dataset-specific features, leading to ineffective verification. For $\mathcal{D}_{f}$, dropping the highest-loss samples causes failure in identifying the independent model. This is probably because dropping highest-loss samples (instead of dropping lowest-loss ones) makes the benign shadow model retain dataset-specific features too similar to those of the victim model. Consequently, the meta-classifier struggles to distinguish independent models that exhibit similar but distinct dataset-specific features, leading to misjudgment.
Table 12: Effects of $\mathcal{D}_{f}$ selection.
$\mathcal{D}_{f}\rightarrow$ Lowest Loss Highest Loss Random